# The Low-Degree Hardness of Finding Large Independent Sets in Sparse Random Hypergraphs

### v2

@@ 4.1 Statistical Threshold
$c_{\gamma}=\frac{1}{r^{r-1}(r-1)\prod_{i}\gamma_{i}},\quad\text{and}\quad f\coloneqq\left(\frac{c_{\bm{\gamma}}\log d}{d}\right)^{1/(r-1)}.$
To prove the upper bound, we will show that
$\mathbb{P}[\alpha_{\bm{\gamma}}(H)\leq(1+\varepsilon)rnf]=1-o(1),\quad\text{as }n\to\infty,$ (8)
@@
$~{}\mathbb{P}[\alpha_{\bm{\gamma}}(H)\geq(1+\varepsilon)rnf]$
$\leq\left(\prod_{i=1}^{r}\binom{n}{\gamma_{i}\,(1+\varepsilon)\,r\,n\,f}\right)\left(1-\frac{d}{n^{r-1}}\right)^{((1+\varepsilon)\,r\,n\,f)^{r}\prod_{i=1}^{r}\gamma_{i}}$
$\leq\left(\prod_{i=1}^{r}\left(\frac{en}{\gamma_{i}\,(1+\varepsilon)\,r\,n\,f}\right)^{\gamma_{i}\,(1+\varepsilon)\,r\,n\,f}\right)\,\exp\left(-(1+\varepsilon)^{r}\,r^{r}\,n\,f^{r}d\prod_{i=1}^{r}\gamma_{i}\right)$
$\leq\exp\left((1+\varepsilon)\,r\,n\,f\left(1-\log r-\log f-\sum_{i=1}^{r}\gamma_{i}\log\gamma_{i}\right)-(1+\varepsilon)^{r}\,r^{r}\,n\,f^{r}d\prod_{i=1}^{r}\gamma_{i}\right)$
$=\exp\left((1+\varepsilon)\,r\,n\,f\left(1-\log r-\log f-\sum_{i=1}^{r}\gamma_{i}\log\gamma_{i}-(1+\varepsilon)^{r-1}r^{r-1}\,c_{\bm{\gamma}}\,\log d\prod_{i=1}^{r}\gamma_{i}\right)\right).$
Note that $-\sum_{i=1}^{r}\gamma_{i}\log\gamma_{i}$ subject to the constraint $\sum_{i}\gamma_{i}=1$ is maximized at $\gamma_{i}=1/r$ for all $i$. With this in hand, plugging in the value for $c_{\bm{\gamma}}$, we may conclude that
$\mathbb{P}[\alpha_{\bm{\gamma}}(H)\geq(1+\varepsilon)rnf]$ $\leq\exp\left((1+\varepsilon)\,r\,n\,f\left(1-\log f-\frac{(1+\varepsilon)^{r-1}\log d}{r-1}\right)\right)$
$=\exp\left((1+\varepsilon)\,r\,n\,f\left(1-\frac{\log c_{\bm{\gamma}}}{r-1}-\frac{((1+\varepsilon)^{r-1}-1)\log d}{r-1}-\frac{\log\log d}{r-1}\right)\right)$
$\leq\exp\left((1+\varepsilon)\,r\,n\,f\left(1-\frac{\log c_{\bm{\gamma}}}{r-1}-\varepsilon\log d-\frac{\log\log d}{r-1}\right)\right)$
@@
$\mathbb{P}[\alpha_{\bm{\gamma}}(H)\geq(1-\varepsilon)rnf]=1-o(1),\quad\text{as }n\to\infty.$ (9)
To this end, we will apply the approach of [perkins2024hardness] for bipartite graphs, which was inspired by the proof of [frieze1990independence] for ordinary graphs. For each $i\in[r]$, denote the vertices of $V_{i}$ as $\{1_{i},\ldots,n_{i}\}$. Let $m=\left(\dfrac{d}{(\log d)^{r}}\right)^{\frac{1}{r-1}}$, $n^{\prime}=n/m$, and let $P_{i,j}=\{((j-1)m+1)_{i},\ldots,(jm)_{i}\}$ for each $i\in[r]$ and $j\in[n^{\prime}]$. We say $I$ is a $\bm{\gamma}$-balanced $P$-independent set if $I$ is $\bm{\gamma}$-balanced and $|X\cap P_{i,j}|\leq 1$ for each $i\in[r]$ and $j\in[n^{\prime}]$. Let $\beta(H)$ denote the size of the largest $\bm{\gamma}$-balanced $P$-independent set in a hypergraph $H\sim\mathcal{H}(r,n,p)$ and let $Z_{k}$ denote the number of such independent sets of size $k$.
@@
$\mathbb{P}\left[|\beta(H)-\overline{\beta}|\geq\lambda\right]\leq 2\exp\left(-\frac{\lambda^{2}m\min_{i\in[r]}\gamma_{i}^{2}}{2rn}\right).$
@@
[…]
  • $v_{l}\in P_{lj_{l}}$ for some $j_{l}>j$ if $l<i$, and
  • $v_{l}\in P_{lj_{l}}$ for some $j_{l}\geq j$ if $l\geq i$.
Alternatively, an edge $(v_{1},\ldots,v_{r})$ is assigned to $E_{ij}$ where $j$ is the minimum “partition number” of a vertex $v\in e$, and $i$ is the minimum value $l$ such that $v_{l}$ is in $P_{lj}$.
For each $1\leq i\leq r$ and $1\leq j\leq n^{\prime}$, we let $Z_{ij}$ denote the outcomes of the edges in $E_{ij}$. Clearly, the variables $Z_{ij}$ are independent. Furthermore, changing the outcomes of the edges in a single $E_{ij}$ can change the value $\beta(\cdot)$ by at most $1/\gamma_{i}$. Therefore, we may apply Lemma 4.2 with $c=1/\min_{i\in[r]}\gamma_{i}$ and $t=rn^{\prime}$ to get
$\mathbb{P}\left[|\beta(H)-\overline{\beta}|\geq\lambda\right]\leq 2\exp\left(-\frac{\lambda^{2}\min_{i\in[r]}\gamma_{i}^{2}}{2rn^{\prime}}\right),$
@@
Let $k=(1-\rho)rnf$ for some $\rho>0$ sufficiently small. Then,
$\mathbb{P}[Z_{k}>0]\geq\exp\left(-\frac{50n\sqrt{r(c_{\gamma}\log d)^{\frac{1}{r-1}}}}{d^{\frac{1-\rho/4}{r-1}}}\right).$
@@
    $u_{\ell}\leq\left(\frac{k\,e}{\ell}\,\frac{k}{n}\,d^{\frac{1-\rho/2}{r-1}}\right)^{\ell}\leq\exp\left(r\,f\,k\,d^{\frac{1-\rho/2}{r-1}}\right)\leq\exp\left(\frac{r^{2}\,n\,\left(c_{\bm{\gamma}}\log d\right)^{2/(r-1)}}{d^{\frac{1+\rho/2}{r-1}}}\right).$
@@
    $\leq\left(r\,f\exp\left(\left(\frac{1-\rho/2}{r-1}\right)\log d\right)\right)^{k}$
    $=\left(\frac{r\,(c_{\gamma}\log d)^{\frac{1}{r-1}}}{d^{\frac{\rho/2}{r-1}}}\right)^{k}\leq 1,$
@@
    $u_{\ell}\leq\left(\frac{e^{2}kn}{(k-\ell)^{2}\,d^{\frac{1-\rho/2}{r-1}}}\right)^{k-\ell}\leq\exp\left(2\sqrt{\frac{kn}{d^{\frac{1-\rho/2}{r-1}}}}\right)=\exp\left(\frac{2n\sqrt{r(c_{\gamma}\log d)^{\frac{1}{r-1}}}}{d^{\frac{1-\rho/4}{r-1}}}\right).$
@@
$\sum_{\ell=0}^{k}u_{\ell}$ $\leq k\left(\exp\left(\frac{r^{2}\,n\,\left(c_{\bm{\gamma}}\log d\right)^{2/(r-1)}}{d^{\frac{1+\rho/2}{r-1}}}\right)+\exp\left(\frac{2n\sqrt{r(c_{\gamma}\log d)^{\frac{1}{r-1}}}}{d^{\frac{1-\rho/4}{r-1}}}\right)\right)$
$\leq\exp\left(\frac{50n\sqrt{r(c_{\gamma}\log d)^{\frac{1}{r-1}}}}{d^{\frac{1-\rho/4}{r-1}}}\right),$
@@ 4.2 Proof of the Achievability Result
In this section, we will construct a degree-$1$ polynomial to prove the first part of Theorem 1.9. In particular, we will construct a polynomial $f\,:\,\{0,1\}^{n^{r}}\to\mathbb{R}^{rn}$ that $(k_{1},\ldots,k_{r},\delta,\xi,0)$-optimizes the maximum $\bm{\gamma}$-balanced independent set problem in $H$ for some $\xi\coloneqq\xi(d,r,\bm{\gamma})\geq 1$ and $\delta=o_{n}(1)$. As mentioned in §1.4, the algorithm we describe is adapted from that of [dhawan2023balanced] for deterministic hypergraphs.
@@
$k_{i}=\gamma_{i}\,r\,n\,(1-\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r.$
If some $k_{i}$ is smaller, we can zero out some coordinates to achieve this. Furthermore, assume $\varepsilon\leq 1/(100r)$ and $i^{\star}=r$.
For each $1\leq i<r$, fix $L_{i}\subseteq V_{i}$ such that $|L_{i}|=k_{i}$ and for vertices $v\in V_{i}$, let $f_{v}(A)=\mathbf{1}\{v\in L_{i}\}$. In particular, $f_{v}$ is a degree-00 polynomial for each $v\in V_{1}\cup\cdots\cup V_{r-1}$. For $v\in V_{r}$, we define
@@
$\frac{k_{r}}{\mu}\leq r\,\gamma_{r}\,\left(\frac{\gamma_{i^{\star}}\log d}{d^{\tilde{\varepsilon}}\,r^{r-1}\,(r-1)\,\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\leq\frac{1}{2},$
@@
$\mathbb{P}\left[\sum_{v\in V_{r}}f_{v}(A)<k_{r}\right]\leq\mathbb{P}\left[\sum_{v\in V_{r}}f_{v}(A)\leq\mu/2\right]=o_{n}(1),$
@@ 4.3 Forbidden Structures
Fix $r\geq 2$, $\varepsilon>0$, and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Additionally, let $i^{\star}=\arg\max_{i}\gamma_{i}$. Let $K\in\mathbb{N}$ be a sufficiently large constant in terms of $r$, $\varepsilon$, and $\bm{\gamma}$. Consider an interpolation path $A^{(0)},\ldots,A^{(T)}$ of length $T=n^{O(r)}$. If $d\coloneqq d(\varepsilon,r,\bm{\gamma},K)$ is sufficiently large, then with probability $1-\exp\left(-\Omega(n)\right)$ there does not exist a sequence of subsets $S_{1},\ldots,S_{K}$ of $V$ satisfying the following:
@@
    for each $k\in[K]$, there exists $t_{k}\in[T]$ such that $S_{k}$ is an independent set in $A^{(t_{k})}$,
@@
    for each $k\in[K]$ and $i\in[r]$, we have $|S_{k}\cap V_{i}|\geq(1+\varepsilon)\,\gamma_{i}\,r\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}$.
@@
    $|S_{k}\setminus\bigcup_{j<k}S_{j}|\in\left[\frac{\tilde{\varepsilon}}{4}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\,\frac{\tilde{\varepsilon}}{2}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\right].$
@@
$\Phi\coloneqq n\left(\frac{\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r-1}\gamma_{i}}\right)^{1/(r-1)},$
@@
$a_{i,k}=\frac{|S_{k}\bigcap V_{i}|}{\Phi},\qquad b_{i,k}=\frac{|(S_{k}\setminus\bigcup_{j<k}S_{j})\cap V_{i}|}{\Phi},\qquad\text{and}\qquad b_{k}=\sum_{i=1}^{k}b_{i,k}.$
Conditions (B2) and (B3) imply that $a_{i,k}\geq(1+\varepsilon)\gamma_{i}r$ and $b_{k}\in[\tilde{\varepsilon}/4,\,\tilde{\varepsilon}/2]$. Then for $c=|\cup_{k}S_{k}|/\Phi$, we have $c\leq\sum_{i}a_{i,k}+(K-1)\tilde{\varepsilon}/2$. As a result of the upper bound in Theorem 1.8, we may assume that $\sum_{i}a_{i,k}\leq(1+\varepsilon)r\,\gamma_{r}^{-1/(r-1)}$. In particular, $c=C(r,\varepsilon,K)$ is a bounded constant independent of $d$. There are at most $n^{2rK}$ choices for $\{a_{i,k}\}$ and $\{b_{i,k}\}$ ($\{b_{k}\}$ can be determined from $\{b_{i,k}\}$). Once these are fixed, the number of tuples $(S_{1},\ldots,S_{K})$ is at most
@@
$\exp\left(\frac{\Phi\,\log d}{r-1}\left(\sum_{i=1}^{r}a_{i,1}+\sum_{k=2}^{K}b_{k}+o(1)\right)\right).$
@@
The goal is to apply the above claim with $x_{i}=\tilde{a}_{i,k}$ and $y_{i}=\tilde{b}_{i,k}$. Note that $\tilde{b}_{i,k}\leq\dfrac{b_{k}}{r\min_{i}\gamma_{i}}\leq\varepsilon/(2r)$ and so we may apply Claim 4.5 with $\varepsilon/2$ to get
@@
$\mathbb{E}[N]$ $\leq n^{2rK}(T+1)^{k}\sup_{\{a_{i,k}\},\{b_{i,k}\}}\exp\left(\frac{\Phi\,\log d}{r-1}\left(\sum_{i=1}^{r}a_{i,1}+\sum_{k=2}^{K}b_{k}+o_{n,d}(1)\right)\right)$
@@
$\leq n^{2rK}(T+1)^{k}\sup_{\{a_{i,k}\},\{b_{i,k}\}}\exp\left(\frac{\Phi\log d}{r-1}\left(r\gamma_{r}(r-1)-\frac{r\varepsilon}{4}\sum_{k=2}^{K}b_{k}+o_{n,d}(1)\right)\right).$
As $b_{k}\geq\tilde{\varepsilon}/4$, for $K$ sufficiently large in terms of $r$ and $\varepsilon$, we have
@@
Next, we show that no independent set in $\mathcal{H}(r,n,p)$ has a large intersection with some fixed set of vertices in all partitions.

### v3

@@ 4.1 Statistical Threshold
$c_{\gamma}=\frac{1}{(r-1)\prod_{i}\gamma_{i}},\quad\text{and}\quad f\coloneqq\left(\frac{c_{\bm{\gamma}}\log d}{d}\right)^{1/(r-1)}.$
Furthermore, we let $q\in\mathbb{N}$ be the smallest integer such that $q\gamma_{i}\in\mathbb{N}$ for each $i\in[r]$. We will assume $d_{0}$ is sufficiently large in terms of $q$. To prove the upper bound, we will show that
$\mathbb{P}[\alpha_{\bm{\gamma}}(H)\leq(1+\varepsilon)nf]=1-o(1),\quad\text{as }n\to\infty,$ (9)
@@
$~\mathbb{P}[\alpha_{\bm{\gamma}}(H)\geq(1+\varepsilon)nf]$
$\leq\left(\prod_{i=1}^{r}\binom{n}{\gamma_{i}\,(1+\varepsilon)\,n\,f}\right)\left(1-\frac{d}{n^{r-1}}\right)^{((1+\varepsilon)\,n\,f)^{r}\prod_{i=1}^{r}\gamma_{i}}$
$\leq\left(\prod_{i=1}^{r}\left(\frac{en}{\gamma_{i}\,(1+\varepsilon)\,n\,f}\right)^{\gamma_{i}\,(1+\varepsilon)\,n\,f}\right)\,\exp\left(-(1+\varepsilon)^{r}\,n\,f^{r}d\prod_{i=1}^{r}\gamma_{i}\right)$
$\leq\exp\left((1+\varepsilon)\,n\,f\left(1-\log f-\sum_{i=1}^{r}\gamma_{i}\log\gamma_{i}\right)-(1+\varepsilon)^{r}\,n\,f^{r}d\prod_{i=1}^{r}\gamma_{i}\right)$
$=\exp\left((1+\varepsilon)\,n\,f\left(1-\log f-\sum_{i=1}^{r}\gamma_{i}\log\gamma_{i}-(1+\varepsilon)^{r-1}c_{\bm{\gamma}}\,\log d\prod_{i=1}^{r}\gamma_{i}\right)\right).$
Note that $\sum_{i=1}^{r}\gamma_{i}\log\gamma_{i}$ subject to the constraint $\sum_{i}\gamma_{i}=1$ is maximized at $\gamma_{i}=1/r$ for all $i$. With this in hand, plugging in the value for $c_{\bm{\gamma}}$, we may conclude that
$\mathbb{P}[\alpha_{\bm{\gamma}}(H)\geq(1+\varepsilon)nf]$ $\leq\exp\left((1+\varepsilon)\,n\,f\left(1+\log r-\log f-\frac{(1+\varepsilon)^{r-1}\log d}{r-1}\right)\right)$
$=\exp\left((1+\varepsilon)\,n\,f\left(1+\log r-\frac{\log c_{\bm{\gamma}}}{r-1}-\frac{((1+\varepsilon)^{r-1}-1)\log d}{r-1}-\frac{\log\log d}{r-1}\right)\right)$
$\leq\exp\left((1+\varepsilon)\,n\,f\left(1+\log r-\frac{\log c_{\bm{\gamma}}}{r-1}-\varepsilon\log d-\frac{\log\log d}{r-1}\right)\right)$
@@
$\mathbb{P}[\alpha_{\bm{\gamma}}(H)\geq(1-\varepsilon)nf]=1-o(1),\quad\text{as }n\to\infty.$ (10)
To this end, we will apply the approach of [PW24] for bipartite graphs, which was inspired by the proof of [Fri90] for ordinary graphs. For each $i\in[r]$, denote the vertices of $V_{i}$ as $\{1_{i},\ldots,n_{i}\}$. Let $m=\left(\dfrac{d}{(\log d)^{r}}\right)^{\frac{1}{r-1}}$, $n^{\prime}=n/m$, and let $P_{i,j}=\{((j-1)m+1)_{i},\ldots,(jm)_{i}\}$ for each $i\in[r]$ and $j\in[n^{\prime}]$. We say $I$ is a $\bm{\gamma}$-balanced $P$-independent set if $I$ is $\bm{\gamma}$-balanced and $|I\cap P_{i,j}|\leq 1$ for each $i\in[r]$ and $j\in[n^{\prime}]$. Let $\beta(H)$ denote the size of the largest $\bm{\gamma}$-balanced $P$-independent set in a hypergraph $H\sim\mathcal{H}(r,n,p)$ and let $Z_{k}$ denote the number of such independent sets of size $k$.
@@
$\mathbb{P}\left[|\beta(H)-\overline{\beta}|\geq\lambda\right]\leq 2\exp\left(-\frac{\lambda^{2}m\,q^{2}}{2rn}\right).$
@@
We will employ the so-called “vertex-based” approach toward applying Theorem 4.2 to problems in random hypergraphs. First, we partition the edges of $K_{r\times n}$ into $rn^{\prime}$ groups $\{E_{ij}\,:\,1\leq i\leq r,\,1\leq j\leq n^{\prime}\}$. We include $(v_{1},\ldots,v_{r})\in E_{ij}$ if and only if the following hold for $1\leq l\leq r$:
  • $v_{l}\in P_{l,j_{l}}$ for some $j_{l}>j$ if $l<i$, and
  • $v_{l}\in P_{l,j_{l}}$ for some $j_{l}\geq j$ if $l\geq i$.
In other words, an edge $e=(v_{1},\ldots,v_{r})$ is assigned to $E_{ij}$ where $j$ is the minimum “partition number” of a vertex $v\in e$, and $i$ is the minimum value $l$ such that $v_{l}$ is in $P_{l,j}$.
For each $1\leq i\leq r$ and $1\leq j\leq n^{\prime}$, we let $Z_{ij}$ denote the outcomes of the edges in $E_{ij}$. Clearly, the variables $Z_{ij}$ are independent. Furthermore, we claim that changing the outcomes of the edges in a single $E_{ij}$ can change the value $\beta(\cdot)$ by at most $q$. We will prove this claim by considering a few cases. To this end, let $\vec{Z}$ denote an arbitrary outcome of the $Z_{ij}$’s, let $H$ denote the hypergraph determined by these outcomes, and let $I$ be the $\bm{\gamma}$-balanced $P$-independent set in $H$ satisfying $|I|=\beta(H)$. Let $\vec{Z}^{\prime}$ be obtained from $\vec{Z}$ by changing the outcomes of the edges in $E_{ij}$ for some $i,j$, let $H^{\prime}$ be the corresponding hypergraph and let $I^{\prime}$ be the largest $\bm{\gamma}$-balanced $P$-independent set in $H^{\prime}$.
  1. (Case 1)
    $I\cap P_{i,j}=\emptyset$. Note that $I$ is still a $\bm{\gamma}$-balanced $P$-independent set in $H^{\prime}$ and so $|I^{\prime}|\geq|I|$. If $I^{\prime}\cap P_{i,j}=\emptyset$, then $I^{\prime}$ is a $\bm{\gamma}$-balanced $P$-independent set in $H$ and so $|I^{\prime}|=|I|$. Suppose $|I^{\prime}\cap P_{i,j}|=1$. Then, there exists a $\bm{\gamma}$-balanced $P$-independent set $J^{\prime}\subseteq I^{\prime}$ in $H^{\prime}$ such that $|J^{\prime}|\geq|I^{\prime}|-q$ and $J^{\prime}\cap P_{i,j}=\emptyset$. As $J^{\prime}$ is also a $\bm{\gamma}$-balanced $P$-independent set in $H$, it follows that $|J^{\prime}|\leq|I|$. In particular, $|I^{\prime}|\leq|I|+q$, as desired.
  2. (Case 2)
    $|I\cap P_{i,j}|=1$. Note that there exists a $\bm{\gamma}$-balanced $P$-independent set $J\subseteq I$ in $H$ such that $|J|\geq|I|-q$ and $J\cap P_{i,j}=\emptyset$. As $J$ is a $\bm{\gamma}$-balanced $P$-independent set in $H^{\prime}$, it follows that $|J|\leq|I^{\prime}|$. In particular, $|I^{\prime}|\geq|I|-q$. If $I^{\prime}\cap P_{i,j}=\emptyset$, then $I^{\prime}$ is a $\bm{\gamma}$-balanced $P$-independent set in $H$, implying $|I^{\prime}|\leq|I|$. If not, an identical argument to the previous case shows that $|I^{\prime}|\leq|I|+q$, as desired.
Therefore, we may apply Lemma 4.2 with $c=q$ and $t=rn^{\prime}$ to get
$\mathbb{P}\left[|\beta(H)-\overline{\beta}|\geq\lambda\right]\leq 2\exp\left(-\frac{\lambda^{2}\,q^{2}}{2rn^{\prime}}\right),$
@@
Let $k=(1-\rho)nf$ for some $\rho>0$ sufficiently small. Then,
$\mathbb{P}[Z_{k}>0]\geq\exp\left(\frac{50n(c_{\gamma}\log d)^{\frac{1}{2(r-1)}}}{d^{\frac{1-\rho/4}{r-1}}}\right).$
@@
    $u_{\ell}\leq\left(\frac{k\,e}{\ell}\,\frac{k}{n}\,d^{\frac{1-\rho/2}{r-1}}\right)^{\ell}\leq\exp\left(f\,k\,d^{\frac{1-\rho/2}{r-1}}\right)\leq\exp\left(\frac{n\,\left(c_{\bm{\gamma}}\log d\right)^{2/(r-1)}}{d^{\frac{1+\rho/2}{r-1}}}\right).$
@@
    $\leq\left(f\exp\left(\left(\frac{1-\rho/2}{r-1}\right)\log d\right)\right)^{k}$
    $=\left(\frac{(c_{\gamma}\log d)^{\frac{1}{r-1}}}{d^{\frac{\rho/2}{r-1}}}\right)^{k}\leq 1,$
@@
    $u_{\ell}\leq\left(\frac{e^{2}kn}{(k-\ell)^{2}\,d^{\frac{1-\rho/2}{r-1}}}\right)^{k-\ell}\leq\exp\left(2\sqrt{\frac{kn}{d^{\frac{1-\rho/2}{r-1}}}}\right)=\exp\left(\frac{2n(c_{\gamma}\log d)^{\frac{1}{2(r-1)}}}{d^{\frac{1-\rho/4}{r-1}}}\right).$
@@
$\sum_{\ell=0}^{k}u_{\ell}$ $\leq k\left(\exp\left(\frac{n\,\left(c_{\bm{\gamma}}\log d\right)^{2/(r-1)}}{d^{\frac{1+\rho/2}{r-1}}}\right)+\exp\left(\frac{2n(c_{\gamma}\log d)^{\frac{1}{2(r-1)}}}{d^{\frac{1-\rho/4}{r-1}}}\right)\right)$
$\leq\exp\left(\frac{50n(c_{\gamma}\log d)^{\frac{1}{2(r-1)}}}{d^{\frac{1-\rho/4}{r-1}}}\right),$
@@ 4.2 Proof of the Achievability Result
In this section, we will construct a degree-$1$ polynomial to prove the first part of Theorem 1.9. In particular, we will construct a polynomial $f\,:\,\{0,1\}^{n^{r}}\to\mathbb{R}^{rn}$ that $(k_{1},\ldots,k_{r},\delta,\xi,0)$-optimizes the maximum $\bm{\gamma}$-balanced independent set problem in $H$ for some $\xi\coloneqq\xi(d,r,\bm{\gamma})\geq 1$ and $\delta=\exp\left(-\Omega(n)\right)$. As mentioned in §1.4, the algorithm we describe is adapted from that of [Dha25] for deterministic hypergraphs.
@@
$k_{i}=\gamma_{i}\,n\,(1-\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r.$
If some $k_{i}$ is smaller, we can zero out some coordinates to achieve this. Furthermore, assume $\varepsilon\leq 1/(100r)$ and $i^{\star}=r$ (recall that $i^{\star}=\arg\max_{i}\gamma_{i}$).
For each $1\leq i<r$, fix $L_{i}\subseteq V_{i}$ such that $|L_{i}|=k_{i}$ and for vertices $v\in V_{i}$, let $f_{v}(A)=\mathbf{1}\{v\in L_{i}\}$. In particular, $f_{v}$ is a degree-0 polynomial for each $v\in V_{1}\cup\cdots\cup V_{r-1}$. For $v\in V_{r}$, we define
@@
$\frac{k_{r}}{\mu}\leq\gamma_{r}\,\left(\frac{\gamma_{i^{\star}}\log d}{d^{\tilde{\varepsilon}}\,(r-1)\,\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\leq\frac{1}{2},$
@@
$\mathbb{P}\left[\sum_{v\in V_{r}}f_{v}(A)<k_{r}\right]\leq\mathbb{P}\left[\sum_{v\in V_{r}}f_{v}(A)\leq\mu/2\right]=\exp\left(-\Omega\left(n\right)\right),$
@@ 4.3 Forbidden Structures
Fix $r\geq 2$, $\varepsilon>0$, and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)\cap\mathbb{Q}$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Additionally, let $i^{\star}=\arg\max_{i}\gamma_{i}$. Let $K\in\mathbb{N}$ be a sufficiently large constant in terms of $r$, $\varepsilon$, and $\bm{\gamma}$. Consider an interpolation path $A^{(0)},\ldots,A^{(T)}$ of length $T=n^{O(r)}$. If $d\coloneqq d(\varepsilon,r,\bm{\gamma},K)$ is sufficiently large, then with probability $1-\exp\left(-\Omega(n)\right)$ there does not exist a sequence of subsets $S_{1},\ldots,S_{K}$ of $V$ satisfying the following:
@@
    for each $k\in[K]$, there exists $t_{k}\in[T]$ such that $S_{k}$ is a $\bm{\gamma}$-balanced independent set in $A^{(t_{k})}$,
@@
    for each $k\in[K]$ and $i\in[r]$, we have $|S_{k}\cap V_{i}|\geq(1+\varepsilon)\,\gamma_{i}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}$.
@@
    $|S_{k}\setminus\bigcup_{j<k}S_{j}|\in\left[\frac{\tilde{\varepsilon}}{4r}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\,\frac{\tilde{\varepsilon}}{2r}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\right].$
@@
$\Phi\coloneqq\frac{n}{r}\left(\frac{\log d}{d\,(r-1)\prod_{i=1}^{r-1}\gamma_{i}}\right)^{1/(r-1)},$
@@
$a_{i,k}=\frac{|S_{k}\bigcap V_{i}|}{\Phi},\qquad b_{i,k}=\frac{|(S_{k}\setminus\bigcup_{j<k}S_{j})\cap V_{i}|}{\Phi},\qquad\text{and}\qquad b_{k}=\sum_{i=1}^{r}b_{i,k}.$
Conditions (B2) and (B3) imply that $a_{i,k}\geq(1+\varepsilon)\gamma_{i}r$ and $b_{k}\in[\tilde{\varepsilon}/4,\,\tilde{\varepsilon}/2]$. Then for $c=|\cup_{k}S_{k}|/\Phi$, we have $c\leq\sum_{i}a_{i,1}(ora_{i,1})+(K-1)\tilde{\varepsilon}/2$. As a result of the upper bound in Theorem 1.8, we may assume that $\sum_{i}a_{i,1}\leq(1+\varepsilon)r\,\gamma_{r}^{-1/(r-1)}$. In particular, $c=C(r,\varepsilon,K)$ is a bounded constant independent of $d$. There are at most $n^{2rK}$ choices for $\{a_{i,k}\}$ and $\{b_{i,k}\}$ ($\{b_{k}\}$ can be determined from $\{b_{i,k}\}$). Once these are fixed, the number of tuples $(S_{1},\ldots,S_{K})$ is at most
@@
$\exp\left(\frac{\Phi\,\log d}{r-1}\left(\sum_{i=1}^{r}a_{i,1}+\sum_{k=2}^{K}b_{k}+o_{d}(1)\right)\right).$
@@
The goal is to apply the above claim with $x_{i}=\tilde{a}_{i,k}$ and $y_{i}=\tilde{b}_{i,k}$. Note that $\tilde{b}_{i,k}\leq\dfrac{b_{k}}{r\min_{i}\gamma_{i}}\leq\varepsilon/(2r)$ and $\tilde{a}_{i,k}\geq 1+\varepsilon$. So we may apply Claim 4.5 with $\varepsilon/2$ to get
@@
$\mathbb{E}[N]$ $\leq n^{2rK}(T+1)^{k}\sup_{\{a_{i,k}\},\{b_{i,k}\}}\exp\left(\frac{\Phi\,\log d}{r-1}\left(\sum_{i=1}^{r}a_{i,1}+\sum_{k=2}^{K}b_{k}+o_{d}(1)\right)\right)$
@@
$\leq n^{2rK}(T+1)^{k}\sup_{\{a_{i,k}\},\{b_{i,k}\}}\exp\left(\frac{\Phi\log d}{r-1}\left(r\gamma_{r}(r-1)-\frac{r\varepsilon}{4}\sum_{k=2}^{K}b_{k}+o_{d}(1)\right)\right).$
As $b_{k}\geq\tilde{\varepsilon}/4$, for $K$ sufficiently large in terms of $r$ and $\tilde{\varepsilon}$, we have
@@
Next, we show that no $\bm{\gamma}$-balanced independent set in $\mathcal{H}(r,n,p)$ has a large intersection with some fixed set of vertices in all partitions.